OpenStudy (anonymous):

What is the circumference of a circle with a diameter of 5 meters? (Use 3.14 for Pi.) 78.5 m 31.4 m 15.7 m 8.14 m

OpenStudy (anonymous):

15.7 m

OpenStudy (anonymous):

2*pie*r 2*3.14*2.5

OpenStudy (hoblos):

r=d/2=5/2=2.5 circumference of a circle = 2πr = 2(3.14)(2.5)=15.7